# Deep Unfolding Network with Spatial Alignment for multi-modal MRI reconstruction

### v1

@@ 3.1.2 The optimization algorithm
Fig. 1: The overall structure of the proposed Deep Unfolding Network with Spatial Alignment (DUN-SA) consists of SAM (Spatial Alignment Module) and RM (Reconstruction Module). The RM is composed of SAAMMLB (Spatial Alignment Augmented Multi-Modal Learning Block), DB (Denoising Block), and DCB (Data Consistency Block). SAM is used to solve spatial alignment task, while RM is for reconstruction task. Specifically, SAAMMLB is used to learn spatial alignment augmented inter-modality prior, DB is used to learn denoising prior, and DCB is used to enforce data consistency constraint.
@@
Update $\phi$: Eq. (5a) describes the cross-modal spatial alignment task. We optimize this by employing the gradient-based algorithm, and the $(t+1)$-th optimization step can be expressed as:
$\phi^{t+1}=\phi^{t}-\alpha\nabla_{\phi}\Psi(x^{t},\mathcal{T}(x_{\text{ref}},\phi^{t})),$ (6)
where $\alpha=\rho\lambda$, $\rho$ is the step size and $\nabla_{\phi}\Psi(x^{t},\mathcal{T}(x_{\text{ref}},\phi^{t}))$ is the gradient of the cross-modal spatial alignment term with respect to the displacement field.
Eq. (5b) represents the image reconstruction task. At this point, we regard the cross-modal spatial alignment term as a regularization term. Given that the HQS method has been proven effective for image inverse problems (Geman and Reynolds, 1992; He et al., 2013), by introducing two auxiliary variables $z$ and $s$, we further transform Eq. (5b) to an unconstrained optimization problem:
@@
where $\beta_{1}$ and $\beta_{2}$ are penalty parameters. As $\beta_{1}$ and $\beta_{2}$ approach infinity, the result of minimizing Eq. (7) will converge to the result of minimizing equation Eq. (5b). The optimization for the $(t+1)$-th iteration can be represented as:
$z^{t+1}$ $=\mathop{\arg\min}_{z}\frac{\beta_{1}}{2}\|{x}^{t}-{z}\|_{2}^{2}+\lambda\Psi(z,\mathcal{T}(x_{\text{ref}},\phi^{t+1})),$ (8a)
$s^{t+1}$ $=\mathop{\arg\min}_{s}\frac{\beta_{2}}{2}\|{x}^{t}-{s}\|_{2}^{2}+\eta R({s}),$ (8b)
$x^{t+1}$ $=\mathop{\arg\min}_{x}\frac{1}{2}\left\|{F_{m}x}-\tilde{{k}}\right\|^{2}+\frac{\beta_{1}}{2}\|{x}-{z}^{t+1}\|_{2}^{2}+\frac{\beta_{2}}{2}\|{x}-{s}^{t+1}\|_{2}^{2}.$ (8c)
Update $z$: Given target modality image $x^{t}$ and aligned reference modality image $\mathcal{T}(x_{\text{ref}},\phi^{t+1})$. We define the proximal operator $\text{prox}_{\frac{\lambda}{\beta_{1}}\Psi(\cdot)}(\cdot)$ such that $\text{prox}_{\frac{\lambda}{\beta_{1}}\Psi(\cdot,\mathcal{T}(x_{\text{ref}},\phi^{t+1}))}({x})=\mathop{\arg\min}_{z}\frac{\beta_{1}}{2}\|{x}-{z}\|_{2}^{2}+\lambda\Psi({z},\mathcal{T}(x_{\text{ref}},\phi^{t+1}))$. Eq. (8a) is then solved by the following equation:
$z^{t+1}=\text{prox}_{\frac{\lambda}{\beta_{1}}\Psi(\cdot,\mathcal{T}(x_{\text{ref}},\phi^{t+1}))}(x^{t}).$ (9)
Update $s$: Given the target modality image $x^{t}$. We also define the proximal operator $\text{prox}_{\frac{\eta}{\beta_{2}}R}(\cdot)$ such that $\text{prox}_{\frac{\eta}{\beta_{2}}R}({x})=\mathop{\arg\min}_{s}\frac{\beta_{2}}{2}\|{x}-{s}\|_{2}^{2}+\eta R({s}).$ The solution to Eq. (8b) can then be expressed as:
$s^{t+1}=\text{prox}_{\frac{\eta}{\beta_{2}}R}(x^{t}).$ (10)
@@
$x^{t+1}$ $=\left(F_{m}^{H}F_{m}+\left(\beta_{1}+\beta_{2}\right)I\right)^{-1}\left(F_{m}^{H}\tilde{k}+\beta_{1}z^{t+1}+\beta_{2}s^{t+1}\right)$
$={F}^{H}\Lambda^{-1}\left(M^{H}\tilde{k}+\beta_{1}{F}z^{t+1}+\beta_{2}{F}s^{t+1}\right),$ (11)
@@ 3.2 Deep Unfolding Network with Spatial Alignment
In Section 3.1, we propose an optimization algorithm for joint alignment-reconstruction model. However, this algorithm has some limitations. Firstly, it requires incorporating the spatial alignment task into the reconstruction process and effectively solving both alignment and reconstruction tasks, making the design of the cross-modal spatial alignment term complicated. Secondly, similar to traditional optimization algorithms, this method requires dozens of iterations to converge, thus increasing time and computational costs significantly. To alleviate these problems, inspired by deep unfolding networks, we unfold the optimization algorithm with the network modules.
@@ 3.2.1 Model Overview
We regard Eq. (6) as the cross-modal spatial alignment process and design the Spatial Alignment Module (SAM) accordingly; we view Eq. (7) as the multi-modal reconstruction task, and propose corresponding Reconstruction Module (RM). Here are detailed introductions to both modules:
Within SAM, we utilize the gradient of the cross-modal spatial alignment term to update the displacement field. However, the gradient is difficult to represent accurately due to the non-linearity and high complexity of the cross-modal spatial alignment term. Therefore, we use a deep network to learn the gradient information and refine the displacement field.
In RM, we iteratively update $z$, $s$, $x$ to achieve high-quality reconstruction results. First, based on Eq. (9), we design the SAAMMLB. By fully leveraging the relationship between corresponding points in aligned reference and target modality images, the SAAMMLB is able to effectively learn inter-modality prior. Further, we consider Eq. (10) as a denoising process and consequently design the DB to learn intra-modality denoising priors. Finally, we design the DCB according to the closed-form solution in Eq. (11). Subsequently, we will provide a detailed explanation of each module.
3.2.2 Spatial Alignment Module (SAM)
@@
Fig. 3: Detailed configurations of SA-Net, Z-Net and S-Net.
In SAM, in order to maintain the form of Eq. (6), considering the complexity of the cross-modal spatial alignment term, we employ a CNN to learn the gradient of the displacement field. This design refines the spatial alignment result by aligning the reference modality with the intermediate reconstruction result of target modality, thus reducing the impact of artifacts in the undersampled image on spatial alignment process.
Throughout the optimization process, the CNN takes the aligned reference modality image from the previous step and the image generated by the reconstruction module, denoted as $({x^{t}},\mathcal{T}(x_{\text{ref}},\phi^{t+1}))$, as input and outputs the corresponding gradient to update $\phi$ as in Eq. (6). It is worth noting that $x^{0}$ is initialized as zero-filled reconstruction, Furthermore, to effectively capture the multi-scale information between images, we employ a CNN architecture known as U-Net (Ronneberger et al., 2015) and name it SA-Net (spatial alignment network). Ultimately, Eq. (6) can be expressed as:
$\phi^{t+1}=\phi^{t}-\alpha\text{SA-Net}({x^{t}},\mathcal{T}(x_{\text{ref}},\phi^{t})).$ (12)
The learnable parameters include the scaling factor $\alpha$, which is initialized to be 1, and the parameters within the SA-Net, which is initialized to be 0, shown in Fig. 2.
3.2.3 Spatial Alignment Augmented Multi-Modal Learning Block (SAAMMLB)
In SAAMMLB, instead of designing a hand-crafted regularization to learn inter-modality prior, we replace the proximal operator in Eq. (9) by a learnable network. Previous works (Xiang et al., 2018; Sun et al., 2019; Xiang et al., 2019; Dar et al., 2020) have demonstrated the effectiveness of U-Net in multi-modal feature extraction. By concatenating aligned images of two modalities along the channel dimension and feeding them into the network, the forward convolution layers will fuse information and features from corresponding positions between channels. In this paper, we adopted a variant of U-Net as the backbone of SAAMMLB, naming it Z-Net (network used to update z). By feeding the target modality image and the aligned reference modality image, denoted as $(x^{t},\mathcal{T}(x_{\text{ref}},\phi^{t+1}))$, into Z-Net, we obtain an image refined with inter-modality prior. Thus, Eq. (9) can be transformed to:
$z^{t+1}=\text{Z-Net}(x^{t},\mathcal{T}(x_{\text{ref}},\phi^{t+1})).$ (13)
However, directly obtaining the refined image might lead to instability in the reconstruction process. Inspired by ResNet structure (He et al., 2016), We carry out this process based on the reconstructed image from the previous step, and further integrating the information learned from the inter-modality prior to stabilize the reconstruction process, resulting in the equation:
$z^{t+1}=x^{t}+\text{Z-Net}(x^{t},\mathcal{T}(x_{\text{ref}},\phi^{t+1})).$ (14)
The structure of Z-Net is illustrated in the Fig. 3. The learnable parameters include those within the Z-Net.
3.2.4 Denoising Block (DB)
In DB, we consider Eq. (10) as a denoising process driven by a denoising prior (Aggarwal et al., 2019; Xin et al., 2022; Yang et al., 2020a). In general, any existing image denoising network can be applied here. Given the simplicity and effectiveness of U-Net in learning the intra-modality prior (Sriram et al., 2020; Xin et al., 2022), we employ U-Net to replace the proximal mapping, naming it S-Net (network used to update s), leading to a transformation of Eq. (10) into:
$u^{t+1}=\text{S-Net}(x^{t})$ (15)
To stabilize both the training and reconstruction processes, we also incorporate the architecture of ResNet, resulting in the following equation:
$u^{t+1}=x^{t}+\text{S-Net}(x^{t})$ (16)
The structure of S-Net is shown in Fig. 3. The learnable parameters include those within the S-Net.
3.2.5 Data Consistency Block (DCB)
In DCB, to maintain the structure of the algorithm, the reconstruction process can be directly expressed by Eq. (11). The closed-form solution can be obtained under the given $(z^{t+1},s^{t+1})$ from SAAMMLB and DB. The learnable parameters include the scaling factors $\beta_{1}$ and $\beta_{2}$, which are both initialized to be 1.
3.3 Network Parameters and Loss Function
Throughout the training of all model parameters, including $\alpha$ and SA-Net in SAM; Z-Net in SAAMMLB; S-Net in DB; and $\beta_{1}$, $\beta_{2}$ in DCB, we optimize them by minimizing the loss function. The training loss for each training pair is defined as the distance between the reconstructed images and the ground truth images. Structural Similarity Loss (SSIM) focuses on image structure, luminance, and contrast, making it well-suited for our reconstruction task. Specifically, the loss function is defined as:
$\mathcal{L}=\sum_{i=1}^{N}\text{SSIM}({x}_{\text{rec},i},{x}_{\text{gt},i})$ (17)

### v2

@@ 3.1.2 The optimization algorithm
Fig. 1: The overall structure of the proposed Deep Unfolding Network with Spatial Alignment (DUN-SA) consists of SAM (Spatial Alignment Module) and RM (Reconstruction Module). The RM is composed of AIPLB (Aligned Inter-modality Prior Learning Block), DB (Denoising Block), and DCB (Data Consistency Block). SAM is used to solve spatial alignment task, while RM is for reconstruction task. Specifically, AIPLB is used to learn aligned inter-modality prior, DB is used to learn denoising prior, and DCB is used to enforce data consistency constraint.
@@
Update $\phi$: Eq. (5a) describes the cross-modal spatial alignment task. We optimize this by employing the gradient-based algorithm, and the $t$-th optimization stage can be expressed as:
$\phi^{t}=\phi^{t-1}-\alpha\nabla_{\phi}\Psi(x^{t-1},\mathcal{T}(x_{\text{ref}},\phi^{t-1})),$ (6)
where $\alpha=\rho\lambda$, $\rho$ is the stage size and $\nabla_{\phi}\Psi(x^{t-1},\mathcal{T}(x_{\text{ref}},\phi^{t-1}))$ is the gradient of the aligned cross-modal prior term with respect to the displacement field.
Eq. (5b) represents the image reconstruction task. At this point, we regard the aligned cross-modal prior term as a regularization term. Given that the HQS algorithm has been proven effective for image inverse problems [9, 10], by introducing two auxiliary variables $z$ and $s$, we further transform Eq. (5b) to an unconstrained optimization problem:
@@
where $\beta_{1}$ and $\beta_{2}$ are penalty parameters. As $\beta_{1}$ and $\beta_{2}$ approach infinity, the result of minimizing Eq. (7) will converge to the result of minimizing equation Eq. (5b). The optimization for the $t$-th iteration can be represented as:
$z^{t}$ $=\mathop{\arg\min}_{z}\frac{\beta_{1}}{2}\|{x}^{t-1}-{z}\|_{2}^{2}+\lambda\Psi(z,\mathcal{T}(x_{\text{ref}},\phi^{t})),$ (8a)
$s^{t}$ $=\mathop{\arg\min}_{s}\frac{\beta_{2}}{2}\|{x}^{t-1}-{s}\|_{2}^{2}+\eta R({s}),$ (8b)
$x^{t}$ $=\mathop{\arg\min}_{x}\frac{1}{2}\left\|{F_{m}x}-\tilde{{k}}\right\|^{2}+\frac{\beta_{1}}{2}\|{x}-{z}^{t}\|_{2}^{2}+\frac{\beta_{2}}{2}\|{x}-{s}^{t}\|_{2}^{2}.$ (8c)
Update $z$: Given target modality image $x^{t-1}$ and aligned reference modality image $\mathcal{T}(x_{\text{ref}},\phi^{t})$. We define the proximal operator $\text{prox}_{\frac{\lambda}{\beta_{1}}\Psi(\cdot)}(\cdot)$ such that $\text{prox}_{\frac{\lambda}{\beta_{1}}\Psi(\cdot,\mathcal{T}(x_{\text{ref}},\phi^{t}))}({x})=\mathop{\arg\min}_{z}\frac{\beta_{1}}{2}\|{x}-{z}\|_{2}^{2}+\lambda\Psi({z},\mathcal{T}(x_{\text{ref}},\phi^{t}))$. Eq. (8a) is then solved by the following equation:
$z^{t}=\text{prox}_{\frac{\lambda}{\beta_{1}}\Psi(\cdot,\mathcal{T}(x_{\text{ref}},\phi^{t}))}(x^{t-1}).$ (9)
Update $s$: Given the target modality image $x^{t-1}$. We also define the proximal operator $\text{prox}_{\frac{\eta}{\beta_{2}}R}(\cdot)$ such that $\text{prox}_{\frac{\eta}{\beta_{2}}R}({x})=\mathop{\arg\min}_{s}\frac{\beta_{2}}{2}\|{x}-{s}\|_{2}^{2}+\eta R({s}).$ The solution to Eq. (8b) can then be expressed as:
$s^{t}=\text{prox}_{\frac{\eta}{\beta_{2}}R}(x^{t-1}).$ (10)
@@
$x^{t}$ $=\left(F_{m}^{H}F_{m}+\left(\beta_{1}+\beta_{2}\right)I\right)^{-1}\left(F_{m}^{H}\tilde{k}+\beta_{1}z^{t}+\beta_{2}s^{t}\right)$
$={F}^{H}\Lambda^{-1}\left(M^{H}\tilde{k}+\beta_{1}{F}z^{t}+\beta_{2}{F}s^{t}\right),$ (11)
@@ 3.2 Deep Unfolding Network with Spatial Alignment
In Section 3.1, we propose an optimization algorithm for the joint alignment-reconstruction model. However, this algorithm has some limitations that affect its efficiency and performance. First, it requires integrating the spatial alignment task with the reconstruction process, which complicates the design of the aligned cross-modal prior term due to the need to effectively solve both tasks simultaneously. Second, similar to traditional optimization algorithms, this method requires dozens of iterations to converge, significantly increasing both the computational cost and the time to solution. To alleviate these problems, inspired by deep unfolding networks, we unfold the optimization algorithm into corresponding network modules.
@@ 3.2.1 Model Overview
The pipeline of the proposed DUN-SA is illustrated in Fig. 1. The entire network structure consists of several key modules. The Spatial Alignment Module (SAM) optimizes the displacement field, which explicitly contribute to the spatial alignment between different modalities. The Reconstruction Module (RM) aims to reconstruct the target modality image. Within RM, there is the Aligned Inter-modality Prior Learning Block (AIPLB), which is responsible for learning inter-modality prior. Additionally, the Denoising Block (DB) focuses on learning intra-modality denoising prior. Lastly, the Data Consistency Block (DCB) integrates both inter- and intra-modality prior to refine the reconstruction result. Notably, SAM and RM have a mutual improvement relationship. Specifically, SAM faces the challenge of aligning a fully-sampled reference image with an under-sampled target image, which may cause alignment errors due to artifacts. By utilizing intermediate reconstructions from RM, which progressively reduces such artifacts, SAM achieves more accurate alignments. Conversely, RM benefits from these improved alignments, which facilitate more precise inter-modality correspondences and result in higher quality reconstructions.
@@
Fig. 3: Detailed configurations of SA-Net, $\text{ProxNet}_{Z}$ and $\text{ProxNet}_{S}$.
SAM: As shown in Fig. 1, the SAM refines the displacement field to improve spatial alignment between modalities. At stage $t$, The network input is composed of the original reference modality image $x_{ref}$ (with the shape of $N_{x}\times N_{y}$), the displacement field from the previous stage $\phi^{t-1}$ (with the shape of $2\times N_{x}\times N_{y}$), and the intermediate reconstructed image $x^{t-1}$ (with the shape of $N_{x}\times N_{y}$). The outputs are the refined displacement field $\phi^{t}$ (with the shape of $2\times N_{x}\times N_{y}$) and the corresponding aligned reference modality image $\mathcal{T}(x_{\text{ref}},\phi^{t})$ (with the shape of $N_{x}\times N_{y}$). Specifically, SAM is built by unfolding the iterative rules in Eq. (6). It starts by warping the original reference modality image using $\phi^{t-1}$, and then concating with $x^{t-1}$ to execute the gradient operator $\nabla_{\phi}\Psi(\cdot)$. Subsequently, the displacement field is updated through gradient descent: $\phi^{t}=\phi^{t-1}-\alpha\nabla_{\phi}\Psi(x^{t-1},\mathcal{T}(x_{\text{ref}},\phi^{t-1}))$. Considering the complexity of the aligned cross-modal prior term, which makes calculating $\nabla_{\phi}\Psi(\cdot)$ difficult, we use SA-Net (Spatial Alignment Network) to replace the gradient operator, resulting in the update $\phi^{t}=\phi^{t-1}-\alpha\text{SA-Net}({x^{t-1}},\mathcal{T}(x_{\text{ref}},\phi^{t-1}))$. Notably, we initialize $\phi^{0}$ with identity transformation based on the shape of $x_{ref}$. The architecture of SAM is shown in Fig. 2. Details on SA-Net are further discussed in Section 3.2.2.
RM: At stage $t$, the inputs to the Reconstruction Module (RM) include the intermediate reconstructed image $x^{t-1}$ (with the shape of $N_{x}\times N_{y}$) and the aligned reference modality image $\mathcal{T}(x_{\text{ref}},\phi^{t-1})$ (with the shape of $N_{x}\times N_{y}$) provided by SAM. The RM aims to enhance the clarity and quality of the image, producing an improved intermediate reconstruction image $x^{t}$ (with the shape of $N_{x}\times N_{y}$) by solving Eq. (5b). Specifically, we utilize HQS algorithm as detailed in Section 3.1.2, to effectively tackle this equation, and design AIPLB, DB and DCB correspondingly.
AIPLB, DB, and DCB: The framework utilizes AIPLB, DB, and DCB to learn inter-modality prior, intra-modality prior, and to refine the reconstruction of the target modality image, respectively. As shown in Fig. 1, at stage $t$, $x^{t-1}$ is concatenated with $\mathcal{T}(x_{\text{ref}},\phi^{t})$ to serve as input into AIPLB for learning inter-modality prior $z^{t}$ (with the shape of $N_{x}\times N_{y}$). Meanwhile, $x^{t-1}$ is fed into DB to learn intra-modality prior $s^{t}$ (with the shape of $N_{x}\times N_{y}$). Subsequently, both types of priors are input into DCB along with under-sampled target modality k-space data $\mathop{\tilde{k}}$ for further refinement. Specifically, AIPLB is constructed by unfolding Eq. (9) and using $\text{ProxNet}_{Z}$ to replace the proximal operator $\text{prox}_{\frac{\lambda}{\beta_{1}}\Psi(\cdot)}(\cdot)$. Furthermore, to stabilize network training, we enable $\text{ProxNet}_{Z}$ to predict the residual by adding a skip connection from the input to the output, thereby obtaining $z^{t}=x^{t-1}+\text{$\text{ProxNet}_{Z}$}(x^{t-1},\mathcal{T}(x_{\text{ref}},\phi^{t}))$. Similarly, for learning the intra-modality prior in DB, we unfold Eq. (10) and use $\text{ProxNet}_{S}$ to replace the proximal operator $\text{prox}_{\frac{\eta}{\beta_{2}}R}(\cdot)$, we also enable $\text{ProxNet}_{S}$ to predict the residual by adding a skip connection from the input to the output. Finally, in DCB, we unfold Eq. (11) and use the obtained inter- and intra-modality prior $z^{t}$ and $s^{t}$, together with under-sampled target modality k-space data $\mathop{\tilde{k}}$, to refine the reconstruction results, obtaining $x^{t}={F}^{H}\Lambda^{-1}\left(M^{H}\mathop{\tilde{k}}+\beta_{1}{F}z^{t}+\beta_{2}{F}s^{t}\right).$ With the k-stage optimization, the proposed DUN-SA can finely reconstruct the target modality image. Further details about network implementation are described in Section 3.2.2.
In this way, the proposed DUN-SA is correspondingly constructed under the guidance of the optimization algorithm, and every network module has its own physical meaning, corresponding to specific iterative step. It is worth noting that we do not need to explicitly calculate the gradient of the aligned cross-modal prior term and regularization term. Instead, the network learns the updates of $\phi$, $z$, $s$ at each stage. Therefore, we can optimize the process of the optimization algorithm by updating the network parameters.
3.2.2 SA-Net, $\text{ProxNet}_{Z}$ and $\text{ProxNet}_{S}$
In SAM, the displacement field $\phi^{t}$ is updated as follows: $\phi^{t}=\phi^{t-1}-\alpha\text{SA-Net}({x^{t-1}},\mathcal{T}(x_{\text{ref}},\phi^{t-1}))$. This update aims to estimate and compensate for the subtle spatial misalignments between the target and reference modalities. To effectively capture the multi-scale information between images, we employ a CNN architecture known as U-Net [24] as the backbone of SA-Net. The learnable parameters include the scaling factor $\alpha$, which is initialized to be 1, and the parameters within the SA-Net, which is initialized to be 0. The SA-Net consists of an encoder and a decoder. As shown in Fig. 3, the encoder consists of four encoding blocks with residual connection, which contain three convolutional layers with 3 × 3 kernels, Batch Normalization layers and Leaky ReLU nonlinearity. Corresponding to the encoder, the decoder also consists of four decoding blocks with residual connection, which contain three convolutional layers with 3 × 3 kernels, Batch Normalization layers and Leaky ReLU nonlinearity. Average pooling and nearest-neighbor upsampling are adopted to change the spatial size of feature maps.
In AIPLB and DB, instead of utilizing hand-crafted regularization to learn inter-modality and intra-modality prior, we substitute the proximal operator with learnable networks, namely $\text{ProxNet}_{Z}$ and $\text{ProxNet}_{S}$ as detailed in Section 3.2.1.
For $\text{ProxNet}_{Z}$, previous studies [35, 29, 36, 3] have underscored the effectiveness of the U-Net architecture in extracting features from multi-modal data. This process involves concatenating aligned images from two modalities along the channel dimension and inputting them into the network, where the forward convolutional layers amalgamate information and features from corresponding channel positions. In this paper, we utilize a variant of U-Net as the backbone architecture for $\text{ProxNet}_{Z}$.
For $\text{ProxNet}_{S}$, we consider Eq. (10) as a denoising process driven by the intra-modality denoising prior [1, 37, 43]. Typically, any established image denoising network could be applied here. Given the simplicity and proven effectiveness of U-Net in learning intra-modality denoising prior [28, 37], we adopted the same structural framework for $\text{ProxNet}_{S}$ as used for $\text{ProxNet}_{Z}$, but with different inputs, as pointed out in Section 3.2.1.
Both $\text{ProxNet}_{Z}$ and $\text{ProxNet}_{S}$ consist of an encoder and a decoder. As shown in Fig. 3, the encoder consists of four encoding blocks, which contain two convolutional layers with 3 × 3 kernels, Instance Normalization layers and Leaky ReLU nonlinearity. Corresponding to the encoder, the decoder also consists of four decoding blocks with residual connection, which contain two convolutional layers with 3 × 3 kernels, Instance Normalization layers and Leaky ReLU nonlinearity. Average pooling and transpose convolutional upsampling are adopted to change the spatial size of feature maps.
3.3 Network parameters and network training
Network parameters are categorized into two sets; $\alpha$ and SA-Net parameters in SAM are denoted as $\omega$, which are utilized to solve the spatial alignment task. Parameters for $\text{ProxNet}_{Z}$ in AIPLB, $\text{ProxNet}_{S}$ in the DB, and $\beta_{1}$, $\beta_{2}$ in DCB, which address the reconstruction task, are collectively denoted as $\theta$. Inspired by [38], in our network training, we use the Structural Similarity Index Measure (SSIM) as our final loss function:
$\mathcal{L}(\omega,\theta)=\frac{1}{N}\sum_{i=1}^{N}(1-\text{SSIM}(x_{\text{rec},i}(\omega,\theta),x_{\text{gt},i}))$